# ShadowNav: Autonomous Global Localization for Lunar Navigation in Darkness

### v2

@@ 5.3.1 Simulation
Figure 11: Simulation absolute localization histogram results for the half survey trajectories. (a) 1% error added (b) 2% error added (c) 3% error added
@@ 5.3.2 Cinder Lakes
Next, we utilized the trajectories collected at Cinder Lakes to validate absolute localization in an analogue Lunar environment. For evaluation, we used a total of three trajectories with two from the South site and one from the north site. One of the south site trajectories, south_vA was connected via artificial points to make it a loop. There is no image data at the locations but the error grows with the distance. This loop utilizes the same image data repeatedly, but allows for a simulation of stability. For the north site trajectory, different combinations of landmarks were utilized with two to three landmarks utilized during each run. The full results of these trajectories with 2% added error are in Table 3. We observed that performance is solid except for the looping trajectory. Additionally, we observed that the filter experienced issues in attaining absolute errors less than $5\text{\,}\mathrm{m}$ along one of the north site trajectories. Figure 12 presents a comparison of the actual trajectory with 2% added error compared with the looped trajectory at 1% and 2% added error. It is shown that there is roughly $400\text{\,}\mathrm{m}$ between craters. With one traverse, the localization can recover from a $10\text{\,}\mathrm{m}$ error. However, we observed that the max error grows while running multiple loops with 2% error and the filter diverges. However, with multiple loops and 1% error, the max error does not grow and the filter maintains stability. Given these results, we can determine that error growth between landmarks should be no more than $5\text{\,}\mathrm{m}$-$6\text{\,}\mathrm{m}$ error and the length of traverses between landmarks can be subsequently developed by assumming 2% relative localization error.
@@
Figure 13: Cinder Lakes crater absolute localization histogram results for single loop trajectories. (a) 1% error added (b) 2% error added (c) 3% error added
@@ 5.4 Benchmark on Flight-Like Computer
Dataset Resolution SGBM max (sec) Detect max (sec)
Cinder Lakes 1295x602 0.48 0.25
Simulation 2039x1425 4.40 0.91
@@
Dataset Num. Particles PF step avg. [seconds] Pf step max [seconds]
Cinder Lakes 50 1.49 2.17
Cinder Lakes 100 2.96 4.29
Cinder Lakes 200 6.00 8.71
Simulation 50 1.71 2.78
Simulation 100 3.41 5.56
Simulation 200 6.96 11.16
In this section, we demonstrate that the proposed ShadowNav approach is capable of being run onboard a flight mission and benchmark our algorithm on a Snapdragon 8155 processor. First, the perception algorithms of SGBMs stereo and discontinuity detection were evaluated and the results are in Table 4. Table 5 shows the results of benchmarking these localization algorithms across a range of several different parameters. The Endurance-A proposal concept [Keane et al., 2022] calls for having the rover stop for approximately ten minutes between drives in order to perform absolute localization. Additionally, long exposure captures are proposed to be captured every $10\text{\,}\mathrm{m}$. Therefore, for a $300\text{\,}\mathrm{m}$ traverse, there could be on the order of 30 iterations to process. In the worst case, a filter using 200 particles would be able to terminate its particle filtering calculation within 8.5 minutes. However, the average case should terminate faster and fulfill the ConOps requirement. The other consideration is to support processing while driving. The Endurance-A concept also proposed 1 minute stops every $10\text{\,}\mathrm{m}$ for the long exposures that this system would use. The perception could likely run during this stop along with potentially the absolute localization processing depending upon other processing needs. Additionally, there is at 20 seconds of driving between long exposures during which a single core could potentially be allocated to perform the localization step. Therefore, we believe that the computation requirements of this system are suitable for use onboard.
@@ 6 Conclusion
Acknowledgments
The authors would like to thank John Elliot, Jeffery Hall, Satish Khanna, Hari Nayar, Issa Nesnas, Curtis Padgett, and Chris Yahnker for their discussions during the development of this work. The research was carried out at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with the National Aeronautics and Space Administration (80NM0018D0004). © 2024. California Institute of Technology. Government sponsorship acknowledged.

### v3

@@ 5.3.1 Simulation
Figure 11: Simulation absolute localization histogram results for the half survey trajectories. (a) 1% error added (b) 2% error added (c) 3% error added.
First, evaluation is performed using data from simulation to determine the best trajectory to observe a landmark crater. This was accomplished by designing three types of trajectories around the same landmarks. The first trajectory type was a straight trajectory which stops every $10\text{\,}\mathrm{m}$ and travels directly past the crater with the nearest approach approximately $10\text{\,}\mathrm{m}$ from the rim. The second trajectory type was a half survey, which was designed so that observation positions occur 180 degrees around the crater with each observation point approximately $7\text{\,}\mathrm{m}$-$13\text{\,}\mathrm{m}$ from the crater rim. The third trajectory type was similar to half survey but had observation points a full 360 degrees around the crater. Table 2 contains the results for all trajectory types with a 2% error growth and Figure 10 is a histogram of final absolute errors of three different trajectories. For each trajectory, we performed Monte Carlo runs of 30 different tests and each test having a 2% relative localization error. From these results, we observe that the half survey and the full survey approaches perform better than the straight trajectory, which is expected. However, performing a full survey versus half survey does not provide additional benefit. The full survey does have some trajectories that perform better but in other trajectories it actually performs worse. Therefore, we believe that performing a half survey provides enough positional variance to constrain the localization.
Figure 13: Cinder Lakes crater absolute localization histogram results for single loop trajectories. (a) 1% error added (b) 2% error added (c) 3% error added.
@@ 5.3.2 Cinder Lakes
Next, we used the trajectories collected at Cinder Lakes to validate absolute localization in an analogue Lunar environment. For evaluation, we used a total of three trajectories with two from the South site and one from the North site. One of the south site trajectories, south_vA was connected via artificial points to make it a loop. There is no image data at the locations but the error grows with the distance. This loop utilizes the same image data repeatedly, but allows for a simulation of stability. For the north site trajectory, different combinations of landmarks were utilized with two to three landmarks per trajectory. The full results of these trajectories with 2% added error are in Table 3. We observe that performance is under $5\text{\,}\mathrm{m}$ average error and less than $10\text{\,}\mathrm{m}$ max error with the exception of the looping trajectory. Additionally, we observe that the filter experienced issues in attaining absolute errors less than $5\text{\,}\mathrm{m}$ along the north site trajectories containing the ND1 and NM12 landmarks. More than 50% of the final errors from these trajectories were greater than $5\text{\,}\mathrm{m}$. Figure 12 presents a comparison of the actual trajectory with 2% added error compared with the looped trajectory at 1% and 2% added error. It is shown that there is roughly $400\text{\,}\mathrm{m}$ between craters. With one traverse, the localization can recover from a $10\text{\,}\mathrm{m}$ error. However, we observed that the max error grows while running multiple loops with 2% error and the filter diverges. However, with multiple loops and 1% error, the max error does not grow and the filter maintains stability. Given these results, we can determine that error growth between landmarks should be no more than $5\text{\,}\mathrm{m}$-$6\text{\,}\mathrm{m}$ error and the length of traverses between landmarks can be subsequently developed by assuming 2% relative localization error.
@@ 5.4 Benchmark on Flight-Like Computer
In this section, we demonstrate that the proposed ShadowNav approach is capable The Endurance-A proposal  [2] calls for the rover to stop for approximately 10 minutes between drives in order to perform absolute localization. Long exposure captures are proposed to be captured every $10\text{\,}\mathrm{m}$. Therefore, for a $300\text{\,}\mathrm{m}$ traverse, there could be on the order of 30 iterations to process. The Endurance-A concept also proposed $1\text{\,}\mathrm{min}$ stops every $10\text{\,}\mathrm{m}$ to capture long exposure images. First, the perception algorithms of SGBMs stereo and discontinuity detection were evaluated and the results are in Table 4.
@@
Dataset Resolution SGBM max (sec) Detect max (sec) Cinder Lakes 1295x602 0.48 0.25 Simulation 2039x1425 4.40 0.91
@@
Dataset Num. Particles PF step avg. [seconds] Pf step max [seconds] Cinder Lakes 50 1.49 2.17 Cinder Lakes 100 2.96 4.29 Cinder Lakes 200 6.00 8.71 Simulation 50 1.71 2.78 Simulation 100 3.41 5.56 Simulation 200 6.96 11.16
Table 5 shows the results of benchmarking these localization algorithms across a range of several different parameters.
@@ 6 CONCLUSION
ACKNOWLEDGEMENTS
The authors would like to thank John Elliott, Jeffery Hall, Satish Khanna, Hari Nayar, Issa Nesnas, Curtis Padgett, and Chris Yahnker for their discussions during the development of this work.